OpenStudy (anonymous):

okay how would i calculate 500 / 8=? * 5

OpenStudy (anonymous):

can you help @damirax

OpenStudy (anonymous):

2

OpenStudy (anonymous):

okay thanks you ^.^

OpenStudy (anonymous):

Thats wrong ....

OpenStudy (anonymous):

and what do you think

OpenStudy (anonymous):

\[\large \sf \frac{500}{8} = 5x\]\[\large \sf x=\frac{500}{8 \times 5}\] I'll let you do the number crunching

OpenStudy (anonymous):

Thats right

OpenStudy (anonymous):

ye